what is the lenght of the wood strip required to frame a photograph of lenght and breadth 34cm and 42cm respectively
1
Expert's answer
2013-01-08T12:12:47-0500
the length of the wood strip equal to perimeterof rectangle with sides 34 and 42 cm
l=P
P=2*34+2*42=68+84=152 cm
l=152 cm - the length of the wood strip